SOLVE FOR X

(3000*10)+(3300*X)+(2500*1.5)+(1000*[86.5-{X+10+1.5}])=279000


Please and Thank you

Answer by nidatanu(4)

(3000*10)+(3300*X)+(2500*1.5)+(1000)=279000
30000 + 3300X + 3750 + (1000 * [86.5 - X + 11.5]) = 279000
33750 + 3300X + (1000 * [98 - X]) = 279000
33750 + 3300X + (98000 - 1000X) = 279000
131750 + 2300X = 279000
2300X = 279000 - 131750
2300X = 147250
X = 147250/2300
X = 64.02